Product of all values of t satisfying


4(23t)−39(22t)+54(2t)+16=0


A) 3


B) 1


C) −4


D) −6

Grade:12

1 Answers

Badiuddin askIITians.ismu Expert
14 years ago

Dear Mayank

let   2t =x  ( x is always positive)

so equation become  4(x3)−39(x2)+54(x)+16=0

   clearly x=2 is its solution

so (x-2) is its one factor

let other 2 roots are a and b

  a+b+2 =39/4    or  a+b =31/4

  a.b.2 =-16/4

  a.b =-2

 solve for a and b 

 a =8  and b=-1/4

 b root is not possible because x is positive

so 2t =2 and 8

 or  t = 1,3

product =1*3 =3

Please feel free to post as many doubts on our discussion forum as you can.
If you find any question Difficult to understand - post it here and we will get you
the answer and detailed  solution very  quickly.

 We are all IITians and here to help you in your IIT JEE preparation.

All the best.
 
Regards,
Askiitians Experts
Badiuddin